Solve using square roots

3x2 = -75


1
Expert's answer
2021-02-24T14:38:30-0500



"3x^2=-75"

"x^2=-25"

"x=\sqrt{-25}=(\sqrt{25}) (\sqrt{-1})"

"x= (5)(i)" or "(-5)(i)"

"x=5i" or "-5i"